OpenStudy (anonymous):

Multiply. Express your answer as a simple fraction or improper fraction. 3 times 3 over 5 equals n

OpenStudy (anonymous):

3 x \[\frac{ 3 }{ 5 }\]

OpenStudy (anonymous):

\[\frac{ 3 }{ 1 }*\frac{ 3 }{ 5 } \to \frac{ 3*3 }{ 1*5 }\] Just solve that to get the improper fraction.

OpenStudy (anonymous):

\[\frac{ 9 }{ 5}\]

OpenStudy (anonymous):

Yep, and if you want to get the simple fraction just see how many 5/5's you can get and then put the remainder as x/5 next to it. (x being the remainder)